Pina Company purchased machinery for $176,400 on January 1, 2017. It is estimated that the machinery will have a useful life of 20 years, salvage value of $14,700, production of 91,400 units, and working hours of 40,900. During 2017, the company uses the machinery for 11,452 hours, and the machinery produces 11,882 units. Compute depreciation under the straight-line, units-of-output, working hours, sum-of-the-years’-digits, and double-declining-balance methods. (Round intermediate calculations to 5 decimal places, e.g. 1.56487 and final answers to 0 decimal places, e.g. 5,125.)

Depreciation
Straight-line $

Units-of-output $

Working hours $

Sum-of-the-years’-digits $

Double-declining-balance $
Answer #1
Depreciation
Straight-line $        8,085
Units-of-output $      21,021
Working hours $      45,276
Sum-of-the-years’-digits $      15,400
Double-declining-balance $      16,170
Depreciation
Straight-line =(176400-14700)/20
Units-of-output =(176400-14700)/91400*11882
Working hours =(176400-14700)/40900*11452
Sum-of-the-years’-digits =(176400-14700)*20/210
Double-declining-balance =2*(176400-14700)/20